OpenStudy (anonymous):

Find the derivative of f(x) = 2 divided by x at x = -2.

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (mathstudent55):

\(f(x) = \dfrac{2}{x}\) Find \(f'(-2)\)

OpenStudy (anonymous):

-1?

OpenStudy (solomonzelman):

it is the same thing as \(f(x)=2\cdot x^{-1}\) apply the power rule

OpenStudy (astrophysics):

So first find the derivative then plug in x = -2, you can use exponent rules to bring the x up so \[f(x)= 2 \times x^{-1}\]

OpenStudy (anonymous):

is the answer -1?

OpenStudy (astrophysics):

Solomon beat me to it :P

OpenStudy (astrophysics):

No, -1 is not right

OpenStudy (mathstudent55):

\(f(x) = 2x^{-1} \) Use the power rule: for \(f(x) = n^n\), \(f'(x) = nx^{n - 1}\)

OpenStudy (solomonzelman):

-1 is f(-2), but NOT f`(-2)

OpenStudy (mathstudent55):

The 2 is just a constant.

OpenStudy (astrophysics):

You have to find the derivative first, what is the derivative of \[f(x) = \frac{ 2 }{ x }\]

OpenStudy (astrophysics):

As others have suggested, use the power rule mathstudent also shows the rule.

OpenStudy (solomonzelman):

the rule here is: \(\large\color{royalblue}{ \displaystyle \frac{ d }{dx}\left( {\rm a}\cdot x^{\rm n}\right) = {\rm a}\cdot\frac{ d }{dx}\left( x^{\rm n}\right)={\rm a}\cdot\left({\rm n}\cdot x^{{\rm n}-1}\right)={\rm an}\cdot x^{n-1}}\)

OpenStudy (solomonzelman):

this is general (for all x except x=0)

OpenStudy (solomonzelman):

now, in your case \(\large\color{royalblue}{ \displaystyle \frac{ d }{dx}\left( {\rm 2}\cdot x^{\rm -1}\right) = ?}\)

OpenStudy (solomonzelman):

basically, I take the constant out, differentiate the x^n part, and multiply the result from the derivative of x^n times the constant.

OpenStudy (solomonzelman):

take a shot pliz

OpenStudy (anonymous):

1/2

OpenStudy (solomonzelman):

I don't think so

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

1

OpenStudy (solomonzelman):

yup, like that

OpenStudy (solomonzelman):

-1/2

OpenStudy (solomonzelman):

why did you say 1?

OpenStudy (solomonzelman):

if you have any questions please ask whatever is troubling you.

OpenStudy (anonymous):

thanks